The wheels of a car have a 24-in. diameter. When the car is being driven so that the wheels make 10 revolutions per second, how far with the car travel in one minute?
Found 2 solutions by askmemath, cleomenius:
Answer by askmemath(368)

Diameter =24inches and since diamter is twice radius Radius = 24/2 =12
10revolutions/sec
In 1 minute there are 60 sec
number of revolutions in 1 min = 10x60 600
Distance covered = Perimeterx Number of revolutions
2xPixRadius x600
2xpix12x600
14400Pi inches
1inch = 2.5 cms approx
14400Pi x 2.5 = 36000Picms
100cms =1metre
36000Pi/100 = 360Pi metres

Answer by cleomenius(959)

I approached the problem as follows:
Circumference = 24 * pi = 75.4 inches
10 revolutions per second 754 inches per second.
754 * 60 = 45240 inches per minute.
3770 feet per minute.
Cleomenius.